1.Find the sum of 60 terms.
5+5*(1/2)+6+6*(1/2)+....


Answer by greenestamps(13200)



The sum is...
5+5(1/2)+6+6(1/2)+...+34+34(1/2)  [60 terms]

=5(1.5)+6(1.5)+...+34(1.5)  [30 terms]

= (5+6+...+34)(1.5)

= ((30)(5+34)/2)(1.5)   [number of terms, times average of first and last, because the sequence is arithmetic]

= 585(1.5) = 877.5